# UniPCB: A Generation-Assisted Detection Framework for PCB Defect Inspection

### 06 May 2026 v1

@@ Abstract
Printed Circuit Board (PCB) defect inspection faces two compounding challenges: scarce and imbalanced defect samples that limit model training, and insufficient feature representation under complex circuit backgrounds. Existing generation methods rely on single-modality conditions with coarse structural control, while detection methods improve architectures without addressing the data bottleneck. To resolve both challenges jointly, we propose a generation-assisted PCB defect inspection framework that integrates controlled defect synthesis with task-specific defect detection. On the generation side, a Multi-modal Condition Generator extracts complementary edge, depth, and text conditions in parallel. A ScaleEncoder then embeds these conditions into the diffusion U-Net at four resolutions, and a Condition Modulation applies FiLM-style spatially-adaptive modulation at each scale, enabling structurally aligned and defect-aware sample synthesis. On the detection side, an Inverted Residual Shift Attention couples self-attention with shift-wise convolution to jointly capture global context and local texture, and a Cross-level Complementary Fusion Block generates pixel-level gates for selective cross-level feature fusion. The synthesized samples directly enrich the detection training set, so that improvements in generation compound with improvements in detection. Extensive experiments on DsPCBSD+ demonstrate that UniPCB achieves mAP@0.5 of 98.0% and mAP@0.5:0.95 of 61.8% on defect detection, surpassing all compared methods, while the generation branch attains an FID of 129.61 and SSIM of 0.619, outperforming existing conditional generation approaches.
@@
PCB defect detection, Diffusion model, Multi-modal data generation.
@@ I Introduction
Printed circuit boards (PCBs) are essential to modern electronics, but complex manufacturing processes can introduce defects such as shorts, opens, and spurious copper that threaten product reliability. Automated optical inspection (AOI) systems typically rely on reference-image comparison and handcrafted feature matching, making them sensitive to lighting variation, board layout changes, and novel defect morphologies, which can lead to high false-negative rates in production environments [69, 14].
Deep learning-based detectors offer substantially greater robustness by learning discriminative representations directly from data. However, their effectiveness depends on the availability of large volumes of labeled defect samples, which is fundamentally difficult to satisfy in PCB manufacturing. Defective boards are rare by design, and defect categories often follow a severely long-tailed distribution. As a result, models trained on limited available data tend to underfit rare defect categories and generalize poorly to unseen board layouts.
To alleviate this data bottleneck, traditional augmentation techniques such as rotation and flipping have been widely adopted, but they fail to capture the complex and variable defect patterns encountered in real production environments. Deep generative models offer a more principled alternative: VAEs [25, 51] and GANs [11] have been applied to defect synthesis, but VAEs tend to produce blurry images due to pixel-wise reconstruction losses, and GANs suffer from training instability and mode collapse. Diffusion models [16, 43, 44, 66, 42, 65] overcome these limitations with superior sample quality and diversity. Nonetheless, their training on natural image datasets limits their effectiveness on industrial images such as PCBs [5, 13], and existing PCB-oriented methods typically rely on a single control condition injected at the initial resolution, providing insufficient structural guidance for the dense, fine-grained circuit patterns unique to PCB images.
@@
Beyond high-quality sample generation, the expressive capability of the detection framework itself is equally critical. DETR [4] leverages end-to-end global modeling to significantly enhance defect perception in complex scenes, but its slow convergence and high computational cost hinder real-time industrial deployment. RT-DETR [68] improves inference speed while maintaining competitive accuracy through optimized attention mechanisms and detection head design, making it a more practical alternative. However, under complex backgrounds or small defects, its feature modeling and multi-scale feature fusion remain insufficient. The challenge of simultaneously optimizing detection accuracy and computational efficiency has been extensively studied in the general detection literature [47], yet achieving this balance for fine-grained PCB defects under dense circuit patterns remains an open problem [50].
@@ II-B PCB Defect Detection
PCB defect detection differs from natural image detection because PCB defects are usually small, sparse, irregularly shaped, and visually similar to normal circuit patterns. Traditional handcrafted or template-matching methods are sensitive to manufacturing variations and therefore generalize poorly across board layouts and imaging conditions. Deep learning based detectors have improved robustness by learning discriminative representations automatically, and recent studies [60, 54, 59, 26, 32, 53, 34] mainly improve PCB defect detection from two perspectives: adaptive feature modeling and multi-scale feature fusion.
Adaptive feature modeling [31, 28, 12] aims to enhance defect-relevant responses while suppressing complex circuit background interference. Liu et al. [31] introduced a multi-head nonlocal transformer module to capture long-range dependencies and focus on defect regions. Guo et al. [12] proposed an Image-to-Instance Contextual Enhancement Component that incorporates global contextual information to reweight proposal features. Dai et al. [6] designed Dynamic Head, which dynamically modulates detection features with scale-aware, spatial-aware, and task-aware attention. These methods demonstrate the importance of adaptive representation learning for subtle defect perception.
@@
Although these methods improve feature representation and localization, they mainly operate at the detector architecture level and typically assume that the training distribution is sufficiently representative. In practical PCB inspection, however, defect samples remain scarce and highly imbalanced, especially for rare defect categories. As a result, architecture-level improvements alone may still be limited by insufficient training data. This motivates our generation-assisted detection pipeline, which combines structured defect synthesis with task-specific feature modeling to address both data scarcity and representation limitations.
@@ V Conclusion and Limitation
In this paper, we present UniPCB, a generation-assisted PCB defect inspection pipeline that integrates controlled defect synthesis with task-specific defect detection. To alleviate data scarcity and class imbalance, we design a Multi-Modal Condition Generator to derive complementary edge, depth, and text conditions from PCB images. Based on these conditions, we introduce a diffusion-based synthesis network with a ScaleEncoder and Conditional Modulation Blocks, which inject multi-resolution control into the denoising process and generate structurally aligned defect samples. To improve defect representation under complex circuit backgrounds, we further develop an Inverted Residual Shift Attention Block for joint local–global feature modeling and a Cross-level Complementary Fusion Block for selective cross-level feature refinement. Experiments on DsPCBSD+ show that the proposed generation-assisted pipeline improves both synthetic sample quality and downstream detection performance over the compared baselines.
Despite these results, UniPCB does not perform end-to-end joint optimization between the generator and detector. Instead, it provides a coordinated pipeline in which the generator supplies structurally controlled defect samples and the detector is designed to exploit the enriched training data. In addition, the current evaluation is mainly conducted on DsPCBSD+-derived datasets, and further validation on PCB data collected from different AOI devices, production lines, and board layouts is needed to assess cross-domain robustness. Exploring detector-feedback-driven generation and broader industrial deployment scenarios remains future work.

### 26 May 2026 v3

@@ Abstract
Accurate and efficient printed circuit board (PCB) defect inspection is critical for ensuring product reliability in automated electronics manufacturing. However, PCB defect inspection faces two compounding challenges: scarce and imbalanced defect samples that limit model training, and insufficient feature representation under complex circuit backgrounds. Existing generation methods often rely on single-modality conditions with coarse structural control, while detection methods mainly improve model architectures without explicitly addressing the data bottleneck. To address these challenges, we propose UniPCB, a generation-assisted PCB defect inspection framework that integrates controlled defect synthesis with task-specific defect detection. On the generation side, a Multi-modal Condition Generator extracts complementary edge, depth, and text conditions in parallel. A ScaleEncoder embeds these conditions into the diffusion U-Net at four resolutions, and a Condition Modulation block applies FiLM-style spatially adaptive modulation at each scale, enabling structurally aligned and defect-aware sample synthesis. On the detection side, an Inverted Residual Shift Attention block couples self-attention with shift-wise convolution to jointly capture global context and local texture, while a Cross-level Complementary Fusion block generates pixel-level gates for selective cross-level feature fusion. The synthesized samples enrich the detection training set, allowing the detector to benefit from more diverse and informative defect observations. Extensive experiments on DsPCBSD+ demonstrate that UniPCB achieves an mAP@0.5 of 98.0% and an mAP@0.5:0.95 of 61.8% on defect detection, outperforming the compared methods. The generation branch also achieves an FID of 129.61 and an SSIM of 0.619, showing improved conditional defect synthesis quality over existing baselines. Code is available at: https://github.com/House-yuyu/UniPCB.
@@
Printed circuit board defect detection, conditional diffusion model, generation-assisted inspection.
@@ I Introduction
Printed circuit boards (PCBs) are essential components in modern electronic products, and their manufacturing quality directly affects the reliability of downstream devices. Complex fabrication processes may introduce defects such as shorts, opens, mouse bites, spurs, and spurious copper, which can degrade electrical connectivity and cause product failure. Automated optical inspection (AOI) systems are widely used for PCB quality control, but traditional reference-image comparison and handcrafted feature matching are sensitive to lighting variations, board layout changes, and novel defect morphologies, often leading to missed detections or false alarms in practical inspection scenarios [1, 2].
Deep learning-based detectors improve robustness by learning discriminative representations directly from data. However, their effectiveness depends on large and well-annotated defect datasets, which are difficult to obtain in PCB manufacturing. First, mature production lines usually have high yield rates, making defective samples naturally rare. Second, PCB layouts vary across products, limiting the transferability of defect samples across board types. Third, defect annotation requires domain expertise and is time-consuming, especially when dense and small defects need to be precisely localized. As a result, PCB defect datasets are often scarce and class-imbalanced, causing detectors to underfit rare defects and generalize poorly to unseen board layouts.
To alleviate this data bottleneck, traditional augmentation techniques such as rotation, flipping, and blurring have been widely adopted. Nevertheless, these operations mainly transform existing samples and cannot sufficiently model the diverse defect morphologies and local structural variations observed in real PCB images. Deep generative models provide a more flexible alternative. VAEs [3, 4] and GANs [5] have been explored for defect synthesis, but VAEs often produce blurry details and GANs may suffer from training instability and mode collapse. Diffusion models [6, 7, 8, 9, 10, 11] offer stronger training stability and better texture modeling ability, making them suitable for fine-grained industrial defect synthesis. However, their training on natural image datasets limits their effectiveness on industrial images such as PCBs [12, 13], and existing PCB-oriented methods typically rely on a single control condition injected at the initial resolution, providing insufficient structural guidance for the dense, fine-grained circuit patterns unique to PCB images.
@@
In addition to data scarcity, the feature representation capability of the detector remains critical. Transformer-based detectors such as DETR [14] and RT-DETR [15] provide strong global modeling ability, while lightweight and efficient detection designs have been widely studied to balance accuracy and computational cost [16, 17]. Recent studies in surface defect and tiny object detection further indicate that fine-grained feature separation, relation reasoning, class knowledge, and regional attention are important for detecting subtle defects under complex backgrounds [18, 19, 20]. Nevertheless, architecture-level improvements alone may still be constrained by insufficient and imbalanced training data [21].
@@ II-B PCB Defect Detection
PCB defect detection differs from natural image detection because PCB defects are usually small, sparse, irregularly shaped, and visually similar to normal circuit patterns. Traditional handcrafted or template-matching methods are sensitive to manufacturing variations and therefore generalize poorly across board layouts and imaging conditions. Deep learning based detectors have improved robustness by learning discriminative representations automatically, and recent PCB defect detection studies [32, 33, 34, 35, 36] mainly improve performance from two perspectives: adaptive feature modeling and multi-scale feature fusion. Beyond PCB-specific scenarios, recent TCSVT studies on industrial defect and anomaly detection have also explored knowledge distillation for rail surface defects [37], differential distillation for unsupervised anomaly localization [38], multi-perspective multimodal fusion for 3D industrial defects [39], and memory-augmented wavelet representations for anomaly detection [40]. These works further indicate that efficient representation learning and robust localization are central issues in industrial visual inspection.
Adaptive feature modeling [41, 42] aims to enhance defect-relevant responses while suppressing complex circuit background interference. Liu et al. [41] introduced a multi-head nonlocal transformer module to capture long-range dependencies and focus on defect regions. Guo et al. [42] proposed an Image-to-Instance Contextual Enhancement Component that incorporates global contextual information to reweight proposal features. Dai et al. [43] designed Dynamic Head, which dynamically modulates detection features with scale-aware, spatial-aware, and task-aware attention. These methods demonstrate the importance of adaptive representation learning for subtle defect perception.
@@
Although these PCB-specific and general industrial inspection methods improve feature representation, localization, and deployment efficiency, they mainly operate at the detector or anomaly modeling level and typically assume that the training distribution is sufficiently representative. In practical PCB inspection, however, defect samples remain scarce and highly imbalanced, especially for rare defect categories. As a result, architecture-level improvements alone may still be limited by insufficient training data. This motivates our generation-assisted detection pipeline, which combines structured defect synthesis with task-specific feature modeling to address both data scarcity and representation limitations.
@@ V Conclusion and Limitation
In this paper, we present UniPCB, a generation-assisted PCB defect inspection pipeline that integrates controlled defect synthesis with task-specific defect detection pipeline. To alleviate data scarcity and class imbalance, we design a Multi-Modal Condition Generator to derive complementary edge, depth, and text conditions from PCB images. Based on these conditions, we introduce a diffusion-based synthesis network with a ScaleEncoder and Conditional Modulation Blocks, which inject multi-resolution control into the denoising process and generate structurally aligned defect samples. To improve defect representation under complex circuit backgrounds, we further develop an Inverted Residual Shift Attention Block for joint local–global feature modeling and a Cross-level Complementary Fusion Block for selective cross-level feature refinement. Experiments on DsPCBSD+ show that the proposed generation-assisted pipeline improves both synthetic sample quality and downstream detection performance over the compared baselines.
Despite these results, UniPCB does not perform end-to-end joint optimization between the generator and detector. Instead, it provides a coordinated pipeline in which the generator supplies structurally controlled defect samples and the detector is designed to exploit the enriched training data. In addition, the current evaluation is mainly conducted on DsPCBSD+-derived datasets, and further validation on PCB data collected from different production lines, AOI devices, and board layouts is needed to assess cross-domain robustness. Exploring detector-feedback-driven generation and broader industrial deployment scenarios remains future work.